# Mitigating Perspective Distortion-induced Shape Ambiguity in Image Crops

### v1

@@ 5 Experiments
Figure 6: Depth visualizations on NYU. We compare the depth predicted by adding KPE encoding to ZoeDepth [4] with the base ZoeDepth model, on $96\times 128$ input crops. We show the depth predictions along with the squared error $\Delta$ w.r.t. to ground truth depth (ranging from dark blue: low to dark red: high, invalid regions are indicated as grey). Our model predicts better depth as evident by lower $\Delta$ (lower intensity red areas). Gains are more prominent in this low resolution setting compared to $384\times 512$ setting (Fig. in supplementary).
5.2 Dense Metric Depth Prediction
Task: Metric depth prediction from a single image is ill-posed due to depth-scale ambiguity [9], so several works [28, 29] focus on predicting depth up to scale. This could be mitigated to some extent by using domain knowledge about specific datasets, as in ZoeDepth [4]. However, this gets worse in the presence of augmentations, e.g. cropping and scaling [39], due to which these augmentations are excluded during training. Here, we investigate whether additional cues from KPE about the crop location & camera intrinsics could help to improve metric depth estimation in the presence of cropping and scaling augmentations.
Method: We adopt the architecture of ZoeDepth [4]. It consists of a transformer-based MiDaS [29] backbone to estimate relative depth, which is passed to a domain specific decoder along with multi-scale image features from the backbone, to predict the metric depth for each dataset separately. The model uses a discretized representation of 3D space (into a volumetric grid, referred to as bins) to first output coarse depth (represented as the center of the bin), which is then refined by an MLP in subsequent layers (referred to as the MetricBins [4] module). The model is trained using the scale-invariant log loss (similar to [3]) with $384\times 512$ images without any cropping and scaling augmentations during training. There are two variants of the model: trained from scratch and initialized from a MiDAS model pre-trained on $12$ datasets (also on $384\times 512$ images).
Method REL $\downarrow$ RMSE $\downarrow$ $\log_{10}$downarrow
$96\times 128$ crops
    ZoeD-X-N 0.182 0.538 0.078
    ZoeD-X-N + KPE dense 0.175 0.516 0.074
    ZoeD-M12-N 0.187 0.536 0.079
    ZoeD-M12-N + KPE dense 0.182 0.516 0.074
$384\times 512$ crops
    ZoeD-X-N 0.090 0.353 0.039
    ZoeD-X-N + KPE dense 0.089 0.350 0.039
    ZoeD-M12-N 0.084 0.337 0.037
    ZoeD-M12-N + KPE dense 0.082 0.329 0.036
Table 2: Dense metric depth prediction on NYU. We test the effectiveness of adding KPE to models trained at different resolutions: $96\times 128$ input with cropping and scaling augmentation, $384\times 512$ input. We observe larger gains in the former, indicating that our encoding helps with cropping and scaling augmentations. The gains are less evident in the $384\times 512$ setting due to strong initialization from MiDaS model pretrained on 12 different datasets. We use the dense variant of KPE here.
Modifications: We use cropping & scaling augmentations during training and operate at 2 sizes: $96\times 128$ $384\times 512$. We do not change the architecture or loss function.
@@ 5.2 Dense Metric Depth Prediction
Incorporating KPE: We explore two design choices to add KPE: (1) with the relative depth & $12\times 16\times 256$ image features predicted by the MiDaS model, before passing it to the MetricBins module, (2) with the downsampled $24\times 32\times 1024$ feature map before being processed by the BeiT [2] module in MiDaS. We interpolate KPE to a resolution of $12\times 16$ for the former and $24\times 32$ for the latter, in case of the dense variant. For the sparse variant, we broadcast it to the required resolution.
@@
Dataset: We use NYU Depth v2 [31], a popular indoor scene dataset, captured over 450+ unique scenes. We consider the same splits as [4], i.e. 23K images for training, 654 images for testing and compare with ZoeDepth [4]. Similar to the training, we evaluate the models at $96\times 128$ and $384\times 512$.
Results: From the results in Tab. 2, we observe that training and evaluating the model at lower resolution crops leads to a considerable drop in performance, even when using the MiDaS initialization. This is expected, as pointed out in several works [39, 29, 6], due to the drawbacks of using cropping and scaling augmentations in these models during training. Next, we consider the impact of adding KPE. We observe larger gains in the lower resolution $96\times 128$ setting with augmentations, indicating that our encoding is indeed useful. We observe little improvement at the $384\times 512$ resolution, which is likely due to the strong initialization provided by the large MiDaS model which is pre-trained on 12 different datasets at the same resolution.
In terms of design choices, we see gains when adding KPE to the downsampled image features before it is passed to the BeiT [2] module. This allows the transformer to incorporate information about the crop location in the camera’s field of view in the multi-scale feature output. We find that adding KPE to the features from the MiDaS backbone along with the predicted relative depth is not very effective. This could be because most of the capacity of the model lies in the MiDaS backbone (327M+ parameters) and the domain-specific decoder at the end may not need additional cues for refining the depth. From Tab. 4, we find the dense variant of the encoding to work better than the sparse variant. This is likely due to the dense prediction nature of this task.
KITTI Dataset
    Cube R-CNN 72.59 44.80 14.68 56.11
    Cube R-CNN + KPE 73.02 47.08 16.34 57.50
nuScenes Dataset
    Cube R-CNN 72.83 50.45 18.76 58.53
    Cube R-CNN + KPE 74.03 51.50 18.27 59.48
KITTI + nuScenes
    Cube R-CNN 70.17 43.83 14.23 54.59
    Cube R-CNN + KPE 73.10 47.51 16.58 57.54
Table 3: 3D Object Detection results. We add KPE to the Cube R-CNN [6] model and compare the AP scores at different IoUs. We see consistent improvements across all metrics in both single dataset and joint dataset training. The gains are more prominent in the mutli-dataset setting since we incorporate camera intrinsic information in KPE. We use the sparse KPE variant here.
KITTI NYU ARCTIC
$\text{AP}_{\text{3D}}^{\text{50}}\uparrow$ $\text{AP}_{\text{3D}}^{\text{70}}\uparrow$ $\text{AP}_{\text{3D}}^{\text{}}\uparrow$ REL $\downarrow$ RMSE $\downarrow$ $\log_{10}$downarrow AAE $\downarrow$ CDev $\downarrow$ Acc $\downarrow$
KPE sparse 47.08 16.34 57.50 0.097 0.388 0.041 5.92 39.39 8.72
KPE dense 45.11 15.16 55.96 0.088 0.353 0.038 6.16 37.71 9.36
Table 4: Ablation on KPE. We compare two variants of KPE: sparse and dense (Sec. 4). We observe that sparse encoding is better for sparse prediction tasks and dense encoding is better for dense prediction tasks.
5.3 3D Object Detection on KITTI and nuScenes
Task: The goal is to predict the 3D bounding box of objects from a single image. Several methods [30, 18, 6] on this task involve predicting 2D bounding box proposals from the features extracted using a backbone, followed by a decoder which uses the features corresponding to the pixels in the box proposal to regress to the 3D bounding box. We investigate whether KPE can augment the box features with additional cues to aid with the 3D regression.
@@
Incorporating KPE: The region proposal network in the backbone predicts 2D box proposals which are used for detecting objects. The positive box proposals with detected objects are then passed to the cube head to regress the 3D bounding box. We incorporate KPE along with the features of the positive 2D box proposals before feeding to the cube head for 3D regression. Similar to Sec. 5.1 and Sec. 5.2, we explore both the sparse and the dense variant of our encoding. For the sparse variant, we broadcast the features to match the feature resolution whereas for the dense variant, we interpolate the features to the desired feature resolution.
Dataset: We experiment on two commonly used datasets in 3D detection: KITTI [13] and nuScenes [7]. We select these datasets since they capture urban outdoor scenes with a wide field of view, a setting in which perspective distortion is more prominent. We use the same splits as [6]: (1) KITTI: 2.8K training images & 329 validation images, (2) nuScenes: 18K training & 1.1K validation images, and focus on the Cars category. We consider two settings: (1) single dataset training on KITTI and nuScenes separately, (2) joint dataset training on both KITTI and nuScenes. Since KPE contains camera information, we hypothesize that it should be effective in joint dataset training as well. In the joint training on KITTI and nuScenes, we compare with the Cube-RCNN variant that does not use virtual depth to disentangle the impact of virtual depth in multi-camera setting.
Metrics: Following prior work [6], we utilize average-precision (AP) as our metric to evaluate the 3D object detection models. The predicted 3D cuboid is matched to the ground truth 3D cuboid by computing a 3D intersection-over-union (IoU). We also report the AP over a range of different IoU thresholds.
Results: We observe consistent improvements in AP at different IoU thresholds in both the single dataset and the joint dataset settings in Tab. 3. The gains are more prominent in the mutli-dataset setting since we incorporate camera intrinsic information into our KPE encoding. From the ablation in Tab. 4, we see that the sparse variant performs better than the dense variant. This is similar to the 3D pose estimation results in Sec. 5.1, which may be due to the fact that both of these tasks involve making a sparse 3D prediction, unlike the depth estimation model in Sec. 5.2.
@@ 6 Conclusion
We explore the ambiguity induced by perspective distortion in image crops. We provide an intuitive understanding of this ambiguity using parallelepipeds and propose an intrinsics-aware positional encoding (KPE) that incorporates the location of the image crop in the field of the view of the camera. KPE is effective on several tasks involving prediction of 3D attributes from a single image, i.e. 3D pose estimation of articulated objects in contact, dense metric depth prediction, and 3D object detection. As future work, we could apply KPE to tasks involving the prediction of fine-grained attributes, e.g. full 3D shape from a single image.
Limitations: Our analysis only considered geometry of projection. It will be interesting to assess if shading or other cues help alleviate the ambiguity. Our KPE encoding requires camera intrinsics which are typically present in image metadata but they may not always be available.
Acknowledgements: We thank Matthew Chang, Shaowei Liu, Anand Bhattad & Kashyap Chitta for feedback on the draft, and David Forsyth for useful discussion. This material is based upon work supported by an NSF CAREER Award (IIS2143873), NASA (80NSSC21K1030), an NVIDIA Academic Hardware Grant, and the NCSA Delta System (supported by NSF OCI 2005572 and the State of Illinois).

### v2

@@ 5.2 Application 2: Dense Metric Depth Prediction [32, 4]
Figure 6: Depth visualizations on NYU. We compare the depth predicted by adding KPE encoding to ZoeDepth [4] with the base ZoeDepth model. We show the depth predictions along with the squared error $\Delta$ w.r.t. to ground truth depth (ranging from dark blue: low to dark red: high, invalid regions are indicated as grey). Our model predicts better depth as evident by lower $\Delta$ (lower intensity red areas).
Modifications: We train and test on 3:4 aspect ratio image crops that are rescaled to $384\times 512$ before being fed into the network. We do not change the architecture or the loss function.
@@
Incorporating KPE: We experiment with both sparse and dense KPE. We explore two design choices for where to add KPE: (1) with the relative depth & $12\times 16\times 256$ image features predicted by the MiDaS model, before passing it to the MetricBins module, (2) with the downsampled $24\times 32\times 1024$ feature map before being processed by the BeiT [2] module in MiDaS. We interpolate KPE to a resolution of $12\times 16$ for the former and $24\times 32$ for the latter, in case of the dense variant. For the sparse variant, we broadcast it to the required resolution.
Dataset: We use NYU Depth v2 [32], a popular indoor scene dataset, containing over 450+ unique scenes. We consider the same splits as [4]: 23K images for training, 654 images for testing and compare against ZoeDepth [4].
Results: Across both settings (pre-training and no pre-training), KPE leads tom improved performance over not using it, leading to relative improvements of 5.4% – 12.5% across the different metrics and settings (Table 3 a and b). We also compare to Cam-Conv [11], a past encoding that has been used for single image depth estimation. While Cam-Conv also encodes the location of pixels in the camera’s field of view, it doesn’t encode them using sinusoidal embeddings that are used in KPE, and leads to worse performance.
In terms of design choices, we see more gains when adding KPE to the downsampled image features before it is passed to the BeiT [2] module in the MiDaS backbone than after the MiDaS backbone (Table 4 a and b). Early infusion allows the transformer to incorporate information about the crop location in the camera’s field of view in the multi-scale feature output. From Tab. 4 c, we find the dense variant of the encoding to work better than the sparse variant. This is likely due to the dense prediction nature of this task.
Table 5: 3D Object Detection (Main Results). We add KPE to the Cube R-CNN [6] model and compare the AP scores at different IoUs. We see consistent improvements across all metrics. KPE helps and is better than other encodings. Green numbers denote relative improvements over not using any encodings.
Cube R-CNN 70.17 43.83 14.23 54.59
Cube R-CNN + Image location (dense) 71.97 45.03 15.16 56.86
Cube R-CNN + Image location (sparse) 72.08 45.93 15.32 56.30
Cube R-CNN + Cam-Conv [11] 72.62 46.51 14.59 56.82
Cube R-CNN + KPE (sparse) 73.10 +4.2% 47.51 +8.4% 16.58 +16% 57.54 +5.4%
Table 6: 3D Object Detection (KPE Design Choices). See Section 5.3 for details.
Sparse vs. Dense KPE
    Cube R-CNN + KPE (dense) 72.74 46.75 15.23 56.78
    Cube R-CNN + KPE (sparse) 73.10 47.51 16.58 57.54
5.3 Application 3: 3D Object Detection on KITTI, nuScenes [6, 13, 7]
Task: The goal is to predict 3D bounding boxes for objects, given a single image.
@@ 5.3 Application 3: 3D Object Detection on KITTI, nuScenes [6, 13, 7]
Incorporating KPE: The region proposal network in the backbone predicts 2D box proposals which are used for detecting objects. The positive box proposals with detected objects are then passed to the cube head to regress the 3D bounding box. We incorporate KPE along with the features of the 2D box proposals before feeding to the cube head for 3D regression. Similar to Sec. 5.1 and Sec. 5.2, we explore both the sparse and the dense variant of our encoding. For the sparse variant, we broadcast the features to match the feature resolution whereas for the dense variant, we interpolate the features to the desired feature resolution.
Dataset: We experiment on two commonly used datasets in 3D detection: KITTI [13] and nuScenes [7]. These datasets capture urban outdoor scenes with a wide field of view, a setting in which perspective distortion is more prominent. We use the same splits as [6]: (1) KITTI: 2.8K training images & 329 validation images, (2) nuScenes: 18K training & 1.1K validation images, and focus on the Cars category. We consider the multi dataset setting where a single model is trained and tested on KITTI and nuScenes. Since KPE contains camera information, we hypothesize that it should be effective in joint dataset training as well. In the joint training on KITTI and nuScenes, we compare with the Cube-RCNN variant that does not use virtual depth to disentangle the impact of virtual depth in multi-camera setting.
Metrics: Following prior work [6], we utilize 3D average-precision (AP3D) as our metric to evaluate the 3D object detection models. The predicted 3D cuboid is matched to the ground truth 3D cuboid by computing a 3D intersection-over-union (IoU). We also report the AP3D over a range of different IoU thresholds.
Results: We observe consistent improvements in AP at different IoU thresholds Tab. 5 with 4.2% – 16% relative improvement over not using any encoding.
Table 6 compares KPE’s encoding of box location in the camera field of view to just encoding the box location in the 2D image. We find KPE’s encoding of the box’s location in the camera’s field of view to be more effective. KPE is also more effective than Cam-Conv [11]. Sparse KPE works better than dense KPE.
@@
6 Discussion
Our experiments provide several insights on three representative 3D from a single image tasks: 3D pose estimation of articulated objects in contact, dense metric depth prediction, 3D object detection, spanning dense & sparse predictions and convolutional & transformer architectures. Specifically, across all settings, we found the use of KPE, sparse or dense, to improve performance. For applications involving sparse output (such as 3D pose for objects), sparse KPE worked better; while for applications involving dense output (such as depth prediction), dense KPE worked better. KPE is quite flexible and can be injected in the network at different layers. The best location to inject KPE is problem and method dependent. KPE has two important aspects: (a) use of sinusoids to encode, and (b) use of crop’s location in the camera field of view rather than the 2D image. Both aspects lead to KPE being better than Cam-Conv [11] that represents locations in camera’s field of view but without sinusoids, and uses sinusoidal encoding of the 2D location in image. At the same time, KPE outperforms task specific prior encodings: Cam-Conv [11] for depth prediction and 2D image location for 3D object detection. Lastly ablations on the task of 3D object pose estimation (Sec. 5.1) reveal that encoding both scale and location of the crop is better than just encoding the location and KPE is more effective than other ways of denoting the location of the object, e.g. via a masked image.
Limitations: Our analysis only considers geometry of projection. It will be interesting to assess if shading or other cues help alleviate the ambiguity. KPE requires camera intrinsics which are typically present in image metadata but may not always be available. It would be useful to investigate if KPE could be used with predicted intrinsics [20] in such cases. Jointly estimating instrinsics using appearance cues [20] in the image crop (e.g. by adding a loss on the intrinsics) could also help with 3D predictions. Other future directions include analyzing the impact of KPE on shallower networks, smaller receptive fields, predicting normals or full 3D shape from a single image.
@@ 7 Conclusion
We explore the ambiguity induced by perspective distortion in image crops. We provide an intuitive understanding of this ambiguity using parallelepipeds and propose an intrinsics-aware positional encoding (KPE) that incorporates the location of the image crop in the field of the view of the camera. Experimental evaluation across three 3D from a single image tasks (3D pose estimation of articulated objects in contact, dense metric depth prediction and 3D object detection) reveals the effectiveness of KPE.
Acknowledgements: We thank Matthew Chang, Shaowei Liu, Anand Bhattad and Kashyap Chitta for feedback on the draft, and David Forsyth for useful discussion. This material is based upon work supported by an NSF CAREER Award (IIS2143873), NASA (80NSSC21K1030), an NVIDIA Academic Hardware Grant, and the NCSA Delta System (supported by NSF OCI 2005572 and the State of Illinois).